An electron is launched with a constant horizontal
Answer #1

a) We need to calculate the time in the horizontal axis.

x = v_o t \rightarrow t = \frac{x}{v_0}= \frac{0.2}{2.0\times 10^{7}}= 1\times 10^{-8} s

in the vertical axis:

y = \frac{at^2}{2}= \frac{qEt^2}{2m} = \frac{1.6\times 10^{-19}(50000)(1\times 10^{-8})^2}{2(9.11\times 10^{-31})}= 0.439 m

b) The magnetic field required to undeflected the electron the net force must be zero:

F_e = F_M

qE = qvB

so the magnetic field is:

B = \frac{E}{v}= 2.5\times 10^{-3} T

the direction of the magnetic force must be in the opposite direction of the electric field so using the right hand rule

the magnetic field must be out of the page.